OpenStudy (anonymous):

An airplane is flying at an altitude of 7400 ft. If the pilot wishes to descend at an angle of 9 with the horizontal, at what horizontal distance from the airport should he begin his flight?

OpenStudy (anonymous):

do you mean 9 degrees with the horizontal??

OpenStudy (anonymous):

It shud b8 degrees

OpenStudy (anonymous):

Think of the whole thing s a right angled triangle. The height is 7400ft. The angle with the horizontal is 9 degrees. which means tan(theta) = vertical height/horizontal distance Sp the horizontal distance = 7400/tan (9 degrees in radians) = 7400/0.15838 = 46723.07 ft